OpenStudy (anonymous):

Find the slope of the line passing through the points (–3, 4) and (2, –1).

OpenStudy (austinl):

\(\displaystyle slope=\frac{rise}{run}=\frac{\Delta y}{\Delta x}=\frac{y_1-y_2}{x_1-x_2}\) Do you think you can solve it from here or do you need further help? :)

OpenStudy (anonymous):

-3/1

OpenStudy (anonymous):

A. 1 B. -1 C. 3-5 D. -5/3

OpenStudy (austinl):

@Sammy224 I would have to disagree. @Cookie_009 try this to solve it for yourself :) \((-3,4)\rightarrow(x_1,y_1)\\ (2,-1)\rightarrow(x_2,y_2)\)

OpenStudy (anonymous):

My bad it wold be d since i minus the 2

OpenStudy (austinl):

\(\displaystyle m=\frac{4-(-1)}{-3-2}=\frac{5}{-5}=~?\)